OpenStudy (kitkat16):

A six-sided die of unknown bias is rolled 20 times, and the number 3 comes up 6 times. In the next three rounds (the die is rolled 20 times in each round), the number 3 comes up 6 times, 5 times, and 7 times. The experimental probability of rolling a 3 is %, which is approximately % more than its theoretical probability. (Round off your answers to the nearest integer.)

OpenStudy (kitkat16):

so 24/80 = 30% So the experimental probability is 30 %

OpenStudy (kitkat16):

The theoretical probability is 1/20+1/20+1/20+1/20=1/5=20%

OpenStudy (kitkat16):

The experimental probability of rolling a 3 is 30%, which is approximately ????% more than its theoretical probability. This isn't right. :(

OpenStudy (kitkat16):

@mathstudent55 can you help me

OpenStudy (kitkat16):

@mathstudent55 do you know how to do this?

OpenStudy (kitkat16):

@Michele_Laino can u help please?

OpenStudy (michele_laino):

here we have to apply the binomial distribution

OpenStudy (michele_laino):

for example the probability to get the number 3, five times, after 20 rolls, is given by the subsequent formula: \[\Large probability = \left( {\begin{array}{*{20}{c}} {20} \\ 5 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^5} \times {\left( {\frac{5}{6}} \right)^{20 - 5}}\] where: \[\Large \left( {\begin{array}{*{20}{c}} {20} \\ 5 \end{array}} \right) = \frac{{20!}}{{5!\left( {20 - 5} \right)!}}\]

OpenStudy (kitkat16):

sorry i had stepped away

OpenStudy (michele_laino):

and that probability has to be compared with the experimental probability, which is: 5/20, as I can read from the text of your exercise

OpenStudy (michele_laino):

so you have to compute this: \[\Large probability = \left( {\begin{array}{*{20}{c}} {20} \\ 5 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^5} \times {\left( {\frac{5}{6}} \right)^{20 - 5}} = ...?\] you can use Windows calculator, for example

OpenStudy (kitkat16):

5/20 would be 25% yes

OpenStudy (kitkat16):

is my answer for experimental correct? 30%

OpenStudy (kitkat16):

or am I confusing the 2

OpenStudy (michele_laino):

the experimental probability is right, since it is 5/20=25%. Nevertheless we have to compute the theoretical probability, which is gioven by the subsequent formula: \[\Large probability = \left( {\begin{array}{*{20}{c}} {20} \\ 5 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^5} \times {\left( {\frac{5}{6}} \right)^{20 - 5}} = ...?\]

OpenStudy (michele_laino):

given*

OpenStudy (michele_laino):

do you know how to compute that quantity?

OpenStudy (kitkat16):

yes its the six side die there are 4 rounds of 20 and the rolls add up to 6+6+5+7= 24 so 24/80 =30% yes

OpenStudy (michele_laino):

that's right! that is the experimental probability

OpenStudy (michele_laino):

now we have to compute the theoretical probability

OpenStudy (kitkat16):

this I am confused because I thought I was doing it right but the number comes out less than it should be

OpenStudy (michele_laino):

no, please the experimental probability, namely 24/80 is right!

OpenStudy (kitkat16):

i just reread the questions and think I was reading it wrong :P

OpenStudy (michele_laino):

the most part of work, is to compute the theoretical probability, using the binomial distribution

OpenStudy (kitkat16):

sorry either having problems with open study or internet. keep getting the uh oh owl

OpenStudy (michele_laino):

the theoretical probability, is given by the sum of the single probabilities, namely: \[\Large \begin{gathered} \left( {\begin{array}{*{20}{c}} {20} \\ 5 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^5} \times {\left( {\frac{5}{6}} \right)^{20 - 5}} + \hfill \\ \hfill \\ + 2 \times \left( {\begin{array}{*{20}{c}} {20} \\ 6 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^6} \times {\left( {\frac{5}{6}} \right)^{20 - 6}} + \hfill \\ \hfill \\ + \left( {\begin{array}{*{20}{c}} {20} \\ 7 \end{array}} \right) \times {\left( {\frac{1}{6}} \right)^7} \times {\left( {\frac{5}{6}} \right)^{20 - 7}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (kitkat16):

5/20=25% which would make the final answer The experimental probability of rolling a 3 is 30%, which is approximately 10% more than its theoretical probability.

OpenStudy (kitkat16):

Is this correct

OpenStudy (kitkat16):

help!!!

OpenStudy (amistre64):

isnt the theoretical probability of rolling a 3 just 1/6?

OpenStudy (kitkat16):

yes that is correct

OpenStudy (amistre64):

then the percent of error is just how far off we are ... 3/10 - 1/6 ---------- 1/6

OpenStudy (kitkat16):

but it is rolled 80 times

OpenStudy (amistre64):

the theorectical probability is 1/6 regardless of the number of times its rolled ..

OpenStudy (kitkat16):

the answer would be 4/5

OpenStudy (amistre64):

is that from an answer key?

OpenStudy (kitkat16):

no I have no answer key

OpenStudy (kitkat16):

im trying to fiqure out how to do it but keep getting more confused

OpenStudy (amistre64):

spose i guess that ill have 10 presents for christmas, and i get 8 what is the percent of my error? 2 out of 8 = 1/4 or 25% is this a valid approach?

OpenStudy (kitkat16):

yes

OpenStudy (kitkat16):

each roll of the die i only have 1/6 to get a 3

OpenStudy (amistre64):

from the google ... Percent error -- take the absolute value of the error divided by the theoretical value, then multiply by 100.

OpenStudy (amistre64):

|3/10 - 1/6| is the absolute difference between observed and theoretic 1/6 is the theoretic

OpenStudy (kitkat16):

is that 2/15

OpenStudy (amistre64):

1.8 - 1 = .8

OpenStudy (amistre64):

\[\frac{\dfrac{3}{10}-\dfrac{1}{6}}{\dfrac{1}{6}}\] \[\frac{6}{6}*\frac{\dfrac{3}{10}-\dfrac{1}{6}}{\dfrac{1}{6}}\] \[\frac{\dfrac{18}{10}-\dfrac{6}{6}}{\dfrac{6}{6}}=1.8-1\]

OpenStudy (kitkat16):

Im confused where you got the 1.8

OpenStudy (amistre64):

why, i wrote it all out ...

OpenStudy (kitkat16):

.8

OpenStudy (amistre64):

tell me where i got 1.8 from

OpenStudy (kitkat16):

im not sure tried to fiqure that out

OpenStudy (amistre64):

wll we start with the setup 3/10 - 1/6 ---------- 1/6 we good here? this make sense?

OpenStudy (kitkat16):

yes ok the same 4/5

OpenStudy (kitkat16):

1.8 =4/5

OpenStudy (amistre64):

i dont like dividing by 1/6, id rather divide by 1 6*1/6 = 6/6 = 1 soo, let multiply the whole thing by 6/6 6(3/10 - 1/6) ---------- 6/6 now since we are dividing by 1, the bottom is irrelevant 6(3/10 - 1/6) distribute 18/10 - 6/6 simplify 1.8 - 1

OpenStudy (kitkat16):

.8

OpenStudy (amistre64):

1.8 is not equal to 4/5 you cant be almost 2 and be less than 1

OpenStudy (amistre64):

1.8 = 1 + 4/5 if thats how you want to see it

OpenStudy (amistre64):

.8 = 4/5 but thats not a percent of error, 80% is a percentage

OpenStudy (kitkat16):

no matter what I do I keep coming up with 4/5. Is the error the differnce like 20%

OpenStudy (amistre64):

no, the error is 4/5, or 8/10 or .80 .... .80 * 100 = 80, for 80%

OpenStudy (amistre64):

another way to approach this is we rolled 23, we expected 80/6 = 13 (23 - 13)/13 = .769, or about 77% which is close to 80% depending on how you work the numbers

OpenStudy (kitkat16):

If it is 80% it doesn't make sense to me because I thought the percentage was less than 305 because of the way the question reads

OpenStudy (kitkat16):

less than 30%

OpenStudy (amistre64):

how do we define percent of error?

OpenStudy (amistre64):

100 * (actual - theory)/ theory do we agree with this?

OpenStudy (kitkat16):

100*(3/10 - 1/6)(1/6)

OpenStudy (amistre64):

whatever that value is, is our percentage of error

OpenStudy (kitkat16):

and that value is 80 correct

OpenStudy (amistre64):

yes, we calculated that already or, if we are to use the rolls instead: 80/6 = 13.3333 and its rather impossible to have .333 of a roll, so 13 rolls are expected. 100 * (24 - 13)/13

OpenStudy (amistre64):

how you work the problem is up to you, so choose one and stick with it :)

OpenStudy (kitkat16):

ok thanks